ckitookthw0ck2poland:

Solve the following proportion. 2/15 = 13/x A. 390 B. 1.73 C. 2.31 D. 97.5

Arieonna:

Given 12/15 = 13/x ( cross- multiply ) 2x = 195 ( divide both sides by 2 ) x = 97.5 → D

Arieonna:

hope this helps

Arieonna:

so therefore D would be your answer